### v3

# Two-center harmonic oscillator basis for Skyrme-DFT calculations (I): formalism and Proof of Principle

Abstract
We present a new method to solve the nuclear density functional theory (DFT) equations using a two-center harmonic oscillator basis, incorporating pairing and Coulomb interactions. The goal is to efficiently determine the fission and fusion configurations in nuclei. The Coulomb exchange terms are evaluated exactly, allowing for a novel approach to neck formation without the Slater approximation, which has been commonly used in space coordinate-based approaches. The new method has been implemented in the code hfodd, enabling direct comparison with standard one-center solutions. This first paper focuses on deriving and implementing a methodology based on stable, precise, and exact applications of harmonic oscillator bases for the two fragments, which can either overlap or be separated by arbitrarily large distances. The implementation is tested on two proof-of-principle examples using light nuclei, specifically 8Be and 24Mg.
I Introduction
After over eighty years, fission is still a fascinating and hot research topic. Heavy atomic nuclei belong to the class of mesoscopic systems that exhibit emergent phenomena, which are difficult to describe using fundamental interactions between nucleons directly. Moreover, contrary to various nuclear properties, which can be explained in terms of a smaller set of valence nucleons, all nucleons are simultaneously involved in fission. Therefore, those phenomena are often described in terms of phenomenological models that do not use nucleonic degrees of freedom, hybrid microscopic-macroscopic models, or fully microscopic models based on nuclear density functional theory (DFT) principles Andreyev et al. (2018); Schmidt and Jurado (2018); Krappe and Pomorski (2012); Schunck and Robledo (2016); Bender et al. (2020); Schunck and Regnier (2022).
From the microscopic point of view, in nuclear fission we find the perfect laboratory to test and improve our knowledge of the quantum many-body systems and phenomena: from fusion to decay, but also for stellar evolution, energy production, or quantum entanglement, those are just a few other research areas that can benefit from the advances in the theoretical approach to fission.
It was recently shown that angular momenta carried by fission fragments are intrinsically related to mass and charge distributions after scission. Even though the experimental data and statistical models seem to corroborate that idea Wilson et al. (2021), recent results showed a strong dependency on the scission configurations Scamps (2022). Moreover, we can find a variety of articles that provide explanatory models taking into account even earlier stages in the fission path. For instance, performing time-dependent calculations and triple angular momentum projection (light and heavy fragments and the relative motion) shows that the pre-scission bending mode can explain the experimental data Bulgac et al. (2022). Alternatively, it was also demonstrated that incorporating shell and deformation effects in the moments of inertia of the fragments leads to similar conclusions Randrup and Vogt (2021). Hence, even though statistical models give a valuable bulk approach to the problem, these may fail where more sophisticated microscopic phenomena are at play. This limitation was recently highlighted in the computational model FREYA Verbeke et al. (2015), which does not consider the proper -distributions Scamps (2024); Zhou et al. (2023), one of the key ingredients needed to obtain the relative spin angle between fragments, directly connected to the experiment.
Apart from the angular momentum distribution, it is well-known that many of the fragment properties –such as mass distributions or excitation energies– are rather sensible to the scission configuration Schunck and Robledo (2016). However, in the existing DFT approaches to fission, relying on the adiabatic approximation, the definition of the scission point is more arbitrary than physical. We need to describe an excited nucleus that is largely deformed and generally triaxial, that can be oriented differently in space. And, most importantly, we need to link that system to the nascent fragments. In other words, it would be desirable to have a framework that considers the initial nucleus internally structured in fragments, even before the journey through the fission path. In this sense, combining the well-tested DFT models with a two-center basis for building the single-particle (s.p.) states appears to be a meaningful choice.
In the last decades, we have seen two different approaches employing the two-center method. The first one relies on extending an external potential (usually Wood-Saxon) into a two-center version Scharnweber et al. (1971); Chandra and Mosel (1978) and solving it on a two-center basis; the harmonic oscillator (HO) being the usual choice. After this, the Schrödinger equation is solved using the Green’s function method Diaz-Torres (2008). Even though the results shown for both light Hagino and Ichikawa (2017); von Oertzen (1996) and heavy systems Mirea (2008) look promising, the Coulomb interaction is not accurately treated. Indeed, in some cases, it is just approximated as a charged spherical distribution, which does not treat overlapping fragments properly. The computational burden associated with obtaining the Green’s function for large basis sets, and the bad solvability conditions when s.p. crossings occur, make this approach optimal for light masses, but cumbersome when describing the evolution of heavy nuclei along the fission path.
The second type of approach has a stronger “molecular flavour” as the main idea is expanding the s.p. wave function into two different centers (usually the HO states), in the same way as molecular orbitals are expanded into atomic orbitals, see, e.g. Ref. von Oertzen et al. (2006). In nuclear structure studies, results exist for the clustering, based on the idea of hybridization and covalent binding Freer et al. (2018). In the realm of nuclear reactions, adding molecular continuum states improves the description of the scattering of weakly-bound nuclei Moschini and Diaz-Torres (2021). Although there is a plethora of applications, in practice, these methods are primarily suitable for light systems.
If we look for an approach more suitable for heavy nuclei such as DFT, the number of references decreases drastically. Since the pioneering work of Berger and Gogny Berger and Gogny (1980) there have been only a few applications of this technology to covariant DFT (CDFT) Geng et al. (2007); Li et al. (2024). However, these aforementioned implementations were restricted to axially deformed bases, which can be a handicap when trying to model the fission paths that are known to pass through triaxial shapes. It appears that only by controlling the deformation and orientations of the molecular-like components, along with the separation between them, we may see the real impact of the collectivity of the initial states in the post-scission configurations.
In the present work, we focus on the underpinnings of the two-center (Cartesian) HO basis method and its implementation suitable for the Skyrme energy density functional. We embedded the method in the newest version of the DFT solver hfodd Dobaczewski et al. (2021); J. Dobaczewski, et al. (2024), where the user now can set the deformation and separation of the bases as input data to perform the Hartree-Fock (HF) or HF-Bogoliubov calculations.
The structure of the paper is as follows. In section II we present the theoretical framework, highlighting how the usual one-center HO (OCHO) self-consistent procedure can be reformulated to be used within the two-center HO (TCHO) basis. In particular, we show how to treat the Coulomb interaction exactly, both in the direct and exchange channels. Appendices A–C present all details of the implementation. In sections III and IV, we discuss the results of simple calculations performed for small systems, namely, for 8Be, where the Coulomb interaction is tested, and for the symmetric fission channel of 24Mg. Both Proof-of-Princple TCHO calculations are compared against the standard OCHO results. In section V, we provide final remarks and discuss potential applications of this newly developed method in future research.
II Theoretical framework
In this section, we introduce the TCHO basis states and describe the procedure to solve the Hartree-Fock equations using that basis. Due to its relevance to fission, we also present the method to determine the Coulomb energy in the direct and exchange channels. All new features were implemented in the code hfodd Dobaczewski et al. (2021); J. Dobaczewski, et al. (2024) and we refer the reader to the first publication of the code Dobaczewski and Dudek (1997a), where the corresponding implementation of the OCHO Cartesian basis was defined.
II.1 The two-center harmonic oscillator basis
We first consider the co-axial case of the TCHO, where the principal axes of the two bases coincide and the bases are shifted by a vector coinciding with one of those principal axes. This restriction allows us to present the method concisely and to build the baseline for the presentation of the most general case of arbitrarily shifted and tilted bases. Note that such a restriction does not preclude triaxial deformations of both bases.
We begin by considering the s.p. wave function, used to compute properties of nuclei in the Hartree-Fock method, expanded in the three-dimensional TCHO basis,
where is the index of a given s.p. state, are the expansion coefficients, is the Cartesian position vector, and represents the vector of the Cartesian HO quantum numbers. For clarity, here and below we omit the isospin indices of wave functions and matrices. In the Cartesian representation, for the -th center (denoted by or ), the wave function is the product of the shifted and deformed one-dimensional HO basis states,
Note that the wave functions of both centres are here represented in the common reference frame, that is, we use shifted wave functions and not shifted reference frames. In Eq. (1), for each centre, summation over vector represents the sum over the HO quantum numbers suitably restricted to lowest HO states as defined in the OCHO case in Ref. Dobaczewski and Dudek (1997b).
The one-dimensional components of are defined as
where , , or and the standard HO constants are defined as . Absorbing the factor in the normalized Hermite polynomials , and using the dimensionless variables defined as
the wave functions (3) take the form:
The OCHO basis is trivially obtained by setting and .
Even though code hfodd takes advantage of simplex symmetry to accelerate calculations Dobaczewski and Dudek (1997a), our goal is to describe the complex motion of the fission fragments, such as bending or wriggling Bulgac et al. (2021). Hence, the system is in general not invariant under the simplex transformation. Nevertheless, the matrix structure of coefficients and of every one-body operator, , has the following generic form,
where superscripts and represent signs of the simplex quantum numbers.
II.2 The generalised eigenvalue problem
Deformed and shifted wave functions (3) corresponding to both centers are no longer mutually orthogonal. Hence, the HF equations represented on the non-orthogonal basis must be rewritten as
that is, we need to solve a generalised eigenvalue problem, detailed in the next section. In equation (7), represents the mean-field Hamiltonian matrix, is the matrix of coefficients defined in Eq. (1), is the diagonal matrix of the s.p. energies, and is the norm overlap matrix, given by
where is the identity matrix and
To solve the HF equations using non-orthogonal bases, several strategies have been applied in the past: In the two-center shell model calculations, the basis was usually orthogonalized through the Gram-Schmidt procedure Ong and Scheid (1975), whereas in the HF calculations, a new orthogonal basis was usually defined by diagonalizing the matrix Berger and Gogny (1980). In our case, as in the molecular-physics implementations, we chose Löwdin’s canonical orthogonalization Mayer (2003) method. In nuclear physics, Löwdin’s orthogonalization has been widely used in the context of the generator coordinate method, which leads to the so-called Griffin-Hill-Wheeler equation Ring and Schuck (1980).
The main idea of the canonical orthogonalization is to solve the HF equations in a subspace where the eigenvalues of the norm overlap matrix (8), here referred to as , which are smaller than a certain cutoff threshold , are removed and a smaller set of orthogonal wave functions is built as
where are the eigenvectors of the norm overlap matrix. Then, the non-rectangular transformation fulfils and allows us to transform the generalised eigenvalue problem into the orthogonal case.
When solving the HFB equations, one must consider both - and - channels. In this case, the HFB matrix in the orthogonal basis can be obtained as follows:
where and stand for the mean field and pairing field, respectively, and is the usual chemical potential, which ensures the correct number of particles in the system. Once the HF (HFB) equations are solved, the single (quasi)-particle wave functions can be obtained in the TCHO basis using the same transformation.
II.3 Matrix elements of the Skyrme mean field
One of the crucial ingredients of the local DFT is the particle density in space, needed to compute the matrix elements of the mean field as well as the energy of the system. The local density of nucleons reads
where is the occupation factor of the -th s.p. state. Hence, using the TCHO basis, it can be expanded as
Not only the local density but also all other quasi-local densities that build the Skyrme functional Engel et al. (1975); Perlińska et al. (2004) have the TCHO form of Eq. (13) and the arguments presented below apply also to them.
One of the principal advantages of using the HO basis in the local DFT, which stems from the particular form of the wave functions (5), is the fact that the local densities are always in the form of the products of polynomials and Gaussian factors. In the TCHO basis, omitting for clarity the spin degrees of freedom, this gives
The choice of the Gauss-Hermite quadrature of an appropriate order then allows for evaluating all integrals of densities exactly, that is, overall excellent stability and resilience to the rounding errors of the Gauss-Hermite quadrature allow for obtaining the results within the machine precision, usually of the order of 10-15 for the double-precision arithmetics.
In particular, let us consider the task of evaluating the matrix elements of a given term associated with the Skyrme interaction Engel et al. (1975); Perlińska et al. (2004). Apart from the density-dependent term, which we discuss later, all mean fields are linear in densities and thus also have the form of the products of polynomials and Gaussian factors
Then, we obtain the space part of the matrix element as
and from Eqs. (2), (5), and (15) we then have
We see that the integrand above contains products of four types of Gaussian factors each corresponding to either center or , that is, 16 possible combinations. However, it is easy to see that only 5 partitions of the products of four Gaussian factors suffice, see Appendix A.
To use the Gauss-Hermite quadrature we need to transform every integral of Eq. (17) into the usual structure as
where is a polynomial, and are the weights and nodes of the quadrature, respectively, and is its order. For that purpose, we combine the exponents to obtain one single Gaussian, which defines the lattice of the quadrature in function of different combinations of center indices.
Owing to the properties of the Gaussians, only five different lattices are enough (see Appendix A for full details). However, considering the structure of densities shown in Eqs. (13) and (14), the number of polynomials to be evaluated is notably higher in comparison with the usual one-center bases. In the most demanding case, when pairing and density-dependent (DD) interactions are considered, up to 20 different polynomials must be computed (see Appendix B).
At this point we note that an analogous implementation can be used for non-co-axial bases, that is, those not only arbitrarily shifted and deformed, but also arbitrarily tilted. Indeed, in this case, the s.p. wave functions (5) depend on dimensionless variables given by
where are the 33 orthogonal rotation matrices for centers A an B. Therefore, for non-co-axial bases, the Gauss-Hermite integration lattices have to be built by diagonalising the quadratic form in the coordinates, which defines the Gaussian factor in Eq. (17), and then determining the polynomial factors on those lattices. The use of non-co-axial bases will be discussed in a forthcoming publication.
Regarding the inclusion of pairing correlations, instead of using the antisymmetric pairing tensor , we use the modified density matrix, defined as:
which allows us to apply the same method to the matrix elements of the - channel Dobaczewski et al. (1984). However, it requires evaluating additional polynomials on the quadrature lattices (see Appendix B), which leads to a substantial increase in computational time. Thus, we have opted not to include pairing correlations in our Proof-of-Principle calculations, as they are primarily intended to demonstrate the capabilities of the method rather than to provide realistic results, which will be the focus in future publications.
II.4 Coulomb interaction in the TCHO basis
When aiming to describe fission, or any reaction involving two nuclei, proper treatment of the Coulomb interaction is crucial. The Coulomb force acts along the whole fission path and it affects the neck formation, the interaction between the pre-fragments and, of course, the evolution of each fragment after scission Bender et al. (2020). The advantage of using the TCHO basis lies in the ability to describe all stages of the fission process by employing different shifts of the bases. Hence, the TCHO method implemented here allows us to describe the effects of the Coulomb interactions not only in each fragment but also between them on the way to and after the scission. This is vital to describe, among other observables, the total kinetic energy of the reaction Schunck and Robledo (2016). Our TCHO implementation of the exact Coulomb exchange effects is of particular importance, as its consequences for neck formation and fragment distributions have never been considered so far.
The Coulomb interaction can be represented as
where indices 1 and 2 pertain to coordinates of two interacting particles, are the 22 diagonal spin matrices, and , , and are the standard spin, isospin, and space exchange operators, respectively. The direct, exchange, and pairing matrix elements of the Coulomb interaction can be effectively treated by exchanging the space indices of the direct term, see Ref. Schunck et al. (2017). Therefore, below we discuss only the space part of the direct term, that is,
To evaluate this matrix element by employing the Gauss-Hermite quadratures again, we use the method introduced by Girod and Grammaticos Girod and Grammaticos (1983) and later implemented in numerous codes Dobaczewski et al. (1996, 2009); Stoitsov et al. (2013); Marević et al. (2022); Zurek et al. (2024). The method relies on approximating the Coulomb potential by a sum of Gaussians,
where at any order , the strengths and widths can be evaluated by simple algebraic expressions. The key point now is that the matrix elements become separable, that is, they are products of matrix elements separately evaluated in each Cartesian direction,
where
The last step to compute this integral by the Gauss-Hermite quadrature is to express the product of Gaussians in terms of a Gaussian of the quadratic form in variables and , which is presented in detail in Appendix C.
From this point on, the calculation of Coulomb energies and mean fields in the direct, exchange, and pairing channels proceed in close analogy to those implemented for any other finite range interaction, such as Yukawa or Gogny force, which was described in Refs. Dobaczewski et al. (2009); Schunck et al. (2017),
Even though for the one-center calculations the number of Gaussians that ensure the precision of the expansion within the nuclear volume is less than 10 Dobaczewski and Dudek (1997a), for two fragments separated by large distances this cannot be enough. This is illustrated in Fig. 1, where we show the Coulomb form factor compared with expansion (23) for different numbers of Gaussians . We see that expansion on Gaussians is sufficiently precise for distances of up to about 200 fm. Since the computation time scales linearly with , higher values can be easily accommodated if needed.
III Proof of Principle I: testing accuracy in the case of
III.1 Ground state energies: one-center vs. two-center calculation
The ultimate goal of the TCHO solver is to describe the fission path in a more accurate and, if possible, easier way. For that, we need to describe the initial nucleus before the reaction starts. Here, as proof for validating the method, we have chosen the case of 8Be. Its well-known cluster structure is ideal for putting the TCHO formalism to the test.
First, we compare the OCHO and TCHO ground-state results for different numbers of shells included in the calculations. For the OCHO calculations, we used a deformed basis adapted to the deformation of the ground state energy: for b we have MeV and MeV. On the other hand, for the two-center method, we chose both bases adapted to the properties of spherical 4He, MeV. Furthermore, as the average nuclear radius of 4He is around 1.90 fm (as obtained from the HF OCHO calculation) we set the two centers of the basis at fm and fm. The Skyrme functional used is UNEDF1 Kortelainen et al. (2012), which will be the one to be used in realistic fission calculations.
In Fig. 2(a) we see that the TCHO energies at the given number of shells are always below those obtained for the OCHO bases. This means that the TCHO basis gives a better variational approximation of the 8Be wave function. We also see that the quadrupole moments shown in Fig. 2(b) converge faster and in a smoother way to the results obtained at high values of . Hence, it appears that the ground state of 8Be can be better described by the TCHO basis than by the OCHO basis.
III.2 Coulomb interaction between particles: a test of the numerical precision.
Fig. 3(a) shows the direct part of the Coulomb interaction calculated in 8Be for different numbers of Gaussians in function of the separation between the bases. As one can see, in the scale of the figure, differences between using expansions on 10 or 30 Gaussians are not noticeable. In the TCHO method, one can nicely see when both fragments become separated, as the Coulomb interaction then follows that of two point-like charges. In our case, this occurs at a separation of fm, which corresponds to the total quadrupole moment of b. From that point on, the electric repulsion decreases as , converging to the value of twice the direct Coulomb interaction of the 4He fragments. As a more precise visualization, panel (b) shows, in the double logarithmic scale, the relative direct Coulomb energy between fragments, that is, with twice the values for the 4He fragments subtracted. One can see that the behaviour is up to 40 fm well reproduced, even for .
The structure of two separated well-defined fragments can also be seen in Fig. 4, where the exchange part of the Coulomb interaction is shown. Again we show the total values (a) and the relative interactions between fragments (b). Once the fragments separate, the relative exchange interaction energy goes rapidly to zero, showing an exponential decrease that can only be appreciated in the logarithmic scale. This exponential behaviour is up to 15 fm well reproduced even for 10 Gaussians, and up to 30 fm for 30 Gaussians.
We note that at the scission point, the exact Coulomb exchange energy slightly overshoots the value characterizing the two fragments. In the forthcoming publications, we will systematically study the impact of the exact treatment of the Coulomb exchange on neck formation and fragment distributions. Indeed, in the region of low density, the Slater approximation of the Coulomb exchange is not justified.
IV Proof of Principle II:
12C+12C
As the second Proof of Principle, we analyze the symmetric fission channel in 24Mg. On the one hand, this is a well-tested reaction in the two-center formalism Berger and Gogny (1980). On the other hand, as Berger and Gogny showed, 6 HO shells are enough to describe the problem, so we can analyze the qualities of the method without the usual DFT computational burden of a heavy nucleus.
IV.1 Description of the 24Mg ground state
To benchmark the results for the ground state of 24Mg, we performed the OCHO calculation using a deformed basis adapted to the ground state quadrupole moment of around 1.1 b. In the direction of the -axis we included 17 whereas in the perpendicular directions 12 HO shells. In the case of the TCHO calculation, we kept the same strategy as in the 8Be test, adapting both bases to the properties of the spherical 12C and including up to 10 HO shells. In both calculations, we included 572 HO states (twice the size of the spherical basis). For the basis cutoff parameter in the norm eigenvalues set to , the TCHO method discarded a few tens of states due to the basis non-orthogonality, see section II.1.
To study the dependence on the center separation, we analyzed different values in the range between 0 and 9 fm, as larger values than those lead to two completely separated fragments. For the sake of completeness, we also varied the cutoff in the norm overlap . In Fig. 5, we show the ground-state energies and quadrupole moments in function of the separation of the basis for different cutoffs . In this case, to avoid numerical instabilities, we used the density-independent Skyrme functional SV Satuła et al. (2011), however, the same behaviour was found for other Skyrme parametrizations.
We found that the best TCHO description of the ground-state energy was given by the separation of 4 fm between centers, which was just 0.75 keV above the OCHO result including up to 17 HO shells. Moreover, from Fig. 5 we learned that a cutoff of is sufficient to include the relevant number of states for an accurate description. Related to the deformation of the ground state, we saw a weak increase of the average value of in function of the separation of the centers. However, these variations represent less than 5%.
Using the optimum TCHO basis we performed a more realistic calculation using the SkM* Bartel et al. (1982) interaction, including the full Coulomb interaction. Some of the relevant quantities are summarized in Table 1, in which we compare the results obtained for the OCHO and TCHO bases. In either case, we did not conserve simplex, signature, or parity symmetry. In the OCHO calculation, we enforced the axial symmetry of the solution. In contrast, in the TCHO calculation, this symmetry was rapidly fixed self-consistently due to the separation of the bases along the -axis. Despite the smaller number of shells included in the TCHO calculation, we observe that there are no significant differences in any of the results, so it appears that the TCHO method captures the ground-state properties of 24Mg perfectly well.
def. OCHO | sph. TCHO | |
Kinetic energy [MeV] | 387.09 | 388.15 |
Sum s.p. energies [MeV] | -492.36 | -492.44 |
Coulomb interaction [MeV] | 27.90 | 27.93 |
Spin-orbit [MeV] | -22.35 | -22.41 |
Skyrme functional [MeV] | -594.47 | -595.47 |
Total g.s. energy (Routhian) [MeV] | -179.48 | -179.38 |
[b] | 1.17 | 1.17 |
[b2] | 0.012 | 0.012 |
This is of crucial importance when studying more complex systems as the TCHO computational time is significantly higher, even for relatively small numbers of HO shells. In Fig. 6, we show the CPU times used by code hfodd to perform 10 iterations in both bases. For the most time-consuming case of the HO shells, the TCHO calculations require about a factor of 14 more CPU time than the OCHO calculations.
IV.2 Symmetric fission of 24Mg
The symmetric fission reaction is mainly driven by an increasing quadrupole deformation until a structure of separate fragments can be identified. One can be tempted to vary the distance between centers as an alternative to increasing the quadrupole deformation. However, we found that the transition between a compact nucleus and two separated fragments was not well-defined and we could not identify any meaningful “basis scission point” where the linear dependency, due to the overlapping basis states, suddenly disappeared. This fact motivated our strategy to describe fission in four different stages: (1) the ground state, (2) small deformations until the appearance of a reasonably populated neck, (3) the region of larger deformations when the neck gets thinner until it breaks, and (4) two independent fragments. For each of these four steps, we fix a different separation of the centers, with increasing constraints on giving continuation in the fission path between stages.
In principle, we could choose the optimum basis separation for each value of the quadrupole moment. However, as we can see in Fig. 5, the energy and quadrupole-moment curves are rather flat for a certain range of separations. This illustrates the fact that for small separations, the TCHO bases generate fairly similar linearly independent basis states. Taking this fact into account, for the aforementioned stages (1-3) we fixed to 4, 6, or 10 fm, respectively. For b, instead of constraining the values of quadrupole moments, we used the TCHO bases separated by the values of increasing from 10 to 20 fm. We show the TCHO results in Fig. 7(a), separately for the cases without Coulomb interaction, with Coulomb direct terms only, and with the full direct+exchange terms included.
Although the no-Coulomb results are not realistic, they test the method’s ability to describe the fissioning system. In effect, we can see that from b on, the variation of energy is invisible and it corresponds exactly to the energy of two spherical 12C fragments obtained with the same basis parameters as we used for the TCHO 24Mg calculation. When the Coulomb interaction is included, we observe a small barrier at b for both direct and direct+exchange curves. Furthermore, we found that the energy difference between the direct and direct+exchange Coulomb curves becomes larger at b. Beyond b the electrostatic repulsion starts behaving like , corresponding to the two point-like charged particles separated by a certain distance. Hence, we can identify the scission point in the region around 8 b. This is illustrated in Fig. 8, where we show the isoscalar densities evaluated at 7.5 and 8.0 b. We see that at 8.0 b, the low-density neck vanishes, giving rise to two independent fragments both having a nearly spherical shape.
Another relevant aspect to be mentioned is that the TCHO results were obtained without any additional constraints, aside from the quadrupole moment. Namely, neither particle density in the neck nor hexadecapole moments were fixed. This makes the TCHO-basis approach simpler, describing the fission reaction more naturally. Moreover, in Fig. 7(b) we show the TCHO results compared with the OCHO ones, both including the full Coulomb interaction. For the OCHO basis, at very high quadrupole moments, the energy was not well reproduced, even when 30 HO shells and additional constraints in higher-order multipole moments were included.
In Fig. 9 we show the 24Mg twelve lowest proton and neutron s.p. energies calculated in the TCHO basis, in function of the quadrupole moment, cf. Ref. Berger and Gogny (1980). To better appreciate the evolution of the levels, we included the values for b computed with fm. This scenario is equivalent to using the OCHO basis, as the frequencies used in both centers are the same. Even though the lines connect the orbitals ordered by their energies, we can see a high degree of crossing for deformations smaller than 5 b. Beyond that point, most of the s.p. states can be easily identified until they become (asymptotically) degenerate, due to the structure of two separate 12C fragments (whose states are represented by the dashed lines). In the case of the proton energies, shown in panel (b), the match is not perfect due to the Coulomb repulsion between fragments, which is still noticeable after scission. Our results are in good agreement with those obtained with the two-center shell model Chandra and Mosel (1978), showing the equivalence between exploring and in such a symmetric case.
Finally, related to the Coulomb interaction we found that both direct and exchange terms present the same behaviour when OCHO or TCHO basis are used. Due to the maximum distance between fragments, given by the largest quadrupole deformation considered, we approximated the form factor by Gaussians. For the direct part, shown in Fig. 10(a), the TCHO and OCHO curves exhibit similar behaviour, although small differences start to be noticeable near the scission point, around 7.5-8.0 b. For the exchange part, Fig. 10(b), due to the scale in which it is represented, the differences appear more clearly, with the TCHO results converging to the doubled exchange energies of the fragments and the OCHO results missing about 50 keV asymptotically.
V Conclusions and future work
In this work, we presented for the first time the 3D two-center harmonic oscillator basis method implemented in the code hfodd, which allows for breaking symmetries such as time-reversal, parity, simplex and axiality. The purpose of this first publication is to show the formalism involved as well as two simple playground systems to test the accuracy of the method. In both 8Be and 24Mg we found that the post-scission configurations can be described perfectly thanks to the possibility of choosing the separation of the centers. We also showed that the electrostatic repulsion is well-reproduced provided the Coulomb form factor is approximated accurately enough.
Related to the pre-scission states, we observed that the TCHO basis method can accurately describe cluster configurations, even for a few HO shells in both bases. However, one must choose a reasonable center separation to describe different deformed states. Fortunately, the range of these separations is rather wide, and the penalty for not choosing the optimum value is not big enough to be worrisome. Comparing the same symmetric fission channel in 24Mg, using OCHO and TCHO bases, we observed that the latter produces the same curves. Considering that the number of states in the TCHO method was rather small and we didn’t include constraints in the neck density or higher-order multipole moments, the two-center method provides a simpler and more natural tool to describe complex phenomena such as fission.
In the upcoming publications, we will explore more complex systems such as asymmetric fission considering odd-mass nuclei and -particle emission in heavy nuclei. In those cases, having different deformations in both bases as well as constraining the fragment particle numbers will be crucial. Furthermore, pairing correlations will be anaylized in such systems to see the effect on the formation of fragments and the interaction between them. Even though the backbone of the method is set, there is still work to do. To analyze properly the generation and evolution of angular momenta in fission fragments, allowing the bases to have different orientations seems necessary. Furthermore, to put the adiabatic approximation to the test, we will need to allow the two HO bases, namely, their separations and orientations, to evolve in time Dobaczewski (2019).
Acknowledgements.
This work was partially supported by the STFC Grant No. ST/W005832/1, ST/P003885/1 and ST/V001035/1. We acknowledge the CSC-IT Center for Science Ltd., Finland, for the allocation of computational resources. This project was partly undertaken on the Viking Cluster, which is a high-performance computing facility provided by the University of York. We are grateful for computational support from the University of York High-Performance Computing service, Viking and the Research Computing team.Appendix A Integration of the mean-field matrix elements and energy density
From Eq. (16) we see that the general matrix element between the HO states belonging to centers is computed as
The problem, as stated in section II, is to evaluate this integral numerically using the Gauss-Hermite quadratures. Considering the structure of the operators, each integral is the sum of three terms (only one off-diagonal block is enough due to hermiticity):
Hence, in Eqs. (27)-(29) there appear five different Gaussians, which translates into five different lattices where the quadratures need to be computed. After several algebraic steps, we then define five different scaled coordinates, shown in Table (2).
4-0 | |
3-1 | |
2-2 | |
1-3 | |
0-4 |
To simplify the expressions (and the computational burden), we can rewrite the product of the two Hermite polynomials involved as a finite sum of another Hermite polynomial in the style of Dobaczewski and Dudek (1997a):
where coefficients can be computed numerically via the Gauss-Hermite quadrature again. Taking this fact and the proper algebraic modifications into account, the general TCHO matrix element reads
where
Coefficients contain hidden factors and
where the exponent depends on the HO constants and bases’ shifts,
The very same strategy can be used to evaluate the energy of the functional, where the integrands are products of pairs densities. Therefore, again we deal with the products of four Gaussians, with different combinations of the HO constants and shifts. However, the lattices to compute the quadratures are the same and the polynomials required are evaluated again at the same points. The only difference is in the fact that we need to perform the sum over all four indices, instead of computing only one cross-term as before.
Appendix B Evaluation of densities
In Appendix A, we saw that five different lattices are needed to perform the numerical integration for the matrix elements or energy of the functional. This is because the four indices define just five different combinations of shifts and scaling, summarized in Table 2. However, from Eq. (32) we see that we need to evaluate polynomials that define densities in the different lattice points of .
In the density-dependent (DD) term, the coupling constants are proportional to , with and being (usually) a non-integer parameter of the interaction Schunck (2019). Therefore, the Gauss-Hermite quadratures are no longer exact because, in general, the integrands are not polynomials but products of polynomials and non-integer powers of polynomials. Since the total densities are fairly smooth functions, quadratures defined for the density-independent terms can still be quite precise. Nevertheless, as we discuss below, in the TCHO implementation, those terms require special treatment.
To determine how many different polynomials must be evaluated, we use the following principles:
Due to we only explicitly compute those for . Hence, three different polynomials in five lattices make a total of 15. In the case of the pairing densities though, this doesn’t apply because left and right wave functions are related, respectively, to the lower and upper components of the quasiparticle wave functions Dobaczewski et al. (1984), giving as a result 20 different polynomials.
The polynomial already fixes two of the four indices of the lattice to be used. As a result, if DD terms are not considered, only 3 out of 5 possible lattices must be used. For example, if we take , only those lattices with in Table 2 are used (which are 4-0,3-1,2-2). Then, their number can be reduced to 9 polynomials, or to 12 when including pairing in the calculation.
To summarize these results, in Table 3 we show the number of different polynomials needed, in function of the complexity of the calculation.
w/o pairing dens. (HF) | w/ pairing dens. (HFB) | |
no DD terms (SV, SV,…) | 9 | 12 |
with DD terms (SkM*,UDF1,…) | 15 | 20 |
When using the DD interactions, the so-called rearrangement term is required in the mean field, whose structure is different from the rest of the terms Dobaczewski and Dudek (1997a):
where to simplify the presentation, we omitted the part related to the spin density. Then, the matrix elements of the rearrangement term read
One last remark related to the numerical evaluation of this integral is in order. In the calculations performed in the TCHO basis, we can encounter situations where the true density (not its polynomial part) is extremely small. This is the case, for instance, when the two centers are separated to describe the fission fragments far away from one another. In that case, the rearrangement term can lead to numerical issues when the power of the interaction is smaller than one, due to the division by extremely small values. To avoid this behaviour, at small densities, the rearrangement term must be set to zero.
Appendix C TCHO matrix elements of the Coulomb interaction
To compute the integrals appearing in (25) we need to transform the expression in such a way that and are separable. Let’s treat only the exponent, recalling:
so that, the total Gaussian function in is simply . Expanding all the terms involved in (37), we have the following quadratic form
where
which can be transformed into a new one of the type
For that purpose, let’s write (38) in matrix notation:
where
then, we can apply a linear transformation to eliminate the cross-term in the coordinates (proportional to ). If we insert the next transformation
into (38) we have,
where is diagonal. In other words, is the unitary matrix that diagonalizes . Then, in terms of its eigenvalues,
where
Then, expression (44) reads as
Going back to equation (25) and taking into account that is orthogonal –as it diagonalizes the symmetric matrix – we can change the variables of integration as follows
If one finally completes the squares and makes the right changes of variables, the matrix elements can be computed numerically with the double quadrature as follows
where
and the new lattice is related to the rotated coordinates as:

### v5

@@
 After this, the Schrödinger equation is solved using the Green’s function method Diaz-Torres (2008).
-Even though the results shown for both light Hagino and Ichikawa (2017); von Oertzen (1996) and heavy systems Mirea (2008) look promising, the Coulomb interaction is not accurately treated.
+Even though the results shown for both light Hagino and Ichikawa (2017); von Oertzen (1996) and heavy systems J.
+and W.
+(1972); Mirea (2008) look promising, the Coulomb interaction is not accurately treated.
 Indeed, in some cases, it is just approximated as a charged spherical distribution, which does not treat overlapping fragments properly.
@@
 Although there is a plethora of applications, in practice, these methods are primarily suitable for light systems.
-If we look for an approach more suitable for heavy nuclei such as DFT, the number of references decreases drastically.
-Since the pioneering work of Berger and Gogny Berger and Gogny (1980) there have been only a few applications of this technology to covariant DFT (CDFT) Geng et al.
+In the pursuit of suitable approaches for heavy nuclei, Density Functional Theory (DFT) stands out, although only a few groups utilized the two-center HO (TCHO) basis expansion.
+Since the pioneering work of Berger and Gogny more than 40 years ago Berger and Gogny (1980), a handful of axial applications for the Gogny functional exist in areas such as mass distributions Dubray et al.
+(2008); Regnier et al.
+(2016, 2019), fission barriers, and spontaneous fission lifetimes Berger et al.
+(1989); Lemaître et al.
+(2018).
+However, the Gogny implementation has never been published, leaving its applicability largely unknown.
+In addition, the Coulomb interaction was approximated by that of point-like charges near the scission point Berger et al.
+(1989), which may be inadequate for accurately describing the angular momentum generation of fission fragments.
+In the realm of Covariant DFT (CDFT), more details are available on the TCHO basis method Geng et al.
 (2007); Li et al.
-(2024).
-However, these aforementioned implementations were restricted to axially deformed bases, which can be a handicap when trying to model the fission paths that are known to pass through triaxial shapes.
-It appears that only by controlling the deformation and orientations of the molecular-like components, along with the separation between them, we may see the real impact of the collectivity of the initial states in the post-scission configurations.
-In the present work, we focus on the underpinnings of the two-center (Cartesian) HO basis method and its implementation suitable for the Skyrme energy density functional.
-We embedded the method in the newest version of the DFT solver hfodd Dobaczewski et al.
+(2024), although this approach also relies on axially deformed co-axial basis states.
+So far no implementation of the general framework of the triaxial, shifted, and non-co-axial TCHO basis expansions exists.
+While this strategy has been explored in molecular physics Maintz et al.
+(2016), and some analytical expressions can be applied to the two-center expansion Nazmitdinov et al.
+(1996), there is no available nuclear DFT code that exploits this formalism.
+In this work, we focus on the fundamentals of the two-center (Cartesian) HO basis method and its implementation suitable for the Skyrme energy density functional.
+We provide detailed information on how to compute the TCHO matrix elements relying on numerical integration instead of transformation of coefficients Robledo (2022) and how to address the generalized eigenvalue problem in the Hartree-Fock (HF) and HF-Bogoliubov (HFB) cases.
+This method has been integrated into the latest version of the DFT solver hfodd Dobaczewski et al.
 (2021); J.
 Dobaczewski, et al.
-(2024), where the user now can set the deformation and separation of the bases as input data to perform the Hartree-Fock (HF) or HF-Bogoliubov calculations.
+(2024), where users can specify the deformation and separation of the bases as input data to perform the HF or HFB calculations.
 The structure of the paper is as follows.
-In section II we present the theoretical framework, highlighting how the usual one-center HO (OCHO) self-consistent procedure can be reformulated to be used within the two-center HO (TCHO) basis.
+In section II we present the theoretical framework, highlighting how the usual Skyrme+Coulomb one-center HO (OCHO) self-consistent procedure can be reformulated to be used within the TCHO basis.
 In particular, we show how to treat the Coulomb interaction exactly, both in the direct and exchange channels.
@@
 where are the 33 orthogonal rotation matrices for centers A an B.
-Therefore, for non-co-axial bases, the Gauss-Hermite integration lattices have to be built by diagonalising the quadratic form in the coordinates, which defines the Gaussian factor in Eq.
-(17), and then determining the polynomial factors on those lattices.
+However, even for non-co-axial bases, the integration lattices remain the same, as the exponent of the common Gaussian is invariant under rotations.
+Thus, the expressions shown within this section and the appendices are still valid.
 The use of non-co-axial bases will be discussed in a forthcoming publication.